OpenStudy (anonymous):

Find an equation for the nth term of the arithmetic sequence. a19 = -58, a21 = -164

OpenStudy (aum):

The nth term of an arithmetic sequence is: \(\Large a_n = a_1 + (n-1)d\)

OpenStudy (anonymous):

the formula is an=a1+(n-1)d

OpenStudy (anonymous):

How do I find a1 though? and would the common difference be -3?

OpenStudy (aum):

Put n = 19 in the above equation. Then put n = 21. Solve the two equations to find \(a_1\) and d.

OpenStudy (anonymous):

is this module 8?

OpenStudy (anonymous):

How do you set that up????

OpenStudy (aum):

\(\Large a_n = a_1 + (n-1)d\) put n = 19 \(\Large a_{19} = a_1 + (19-1)d = -58\) \(\Large a_{19} = a_1 + 18d = -58\) Try the same with n = 21

OpenStudy (anonymous):

so I end up with a19=a1 + 18d = -58 and a21 + 20d = -64 now what?

OpenStudy (anonymous):

a21 = a1 + 20d + -64*

OpenStudy (anonymous):

= -64

OpenStudy (aum):

a1 + 18d = -58 a1 + 20d = -164 (-164 not -64 and RHS is a1 not a21) Subtract first from second: 2d = -164 + 58 = -106 d = -106/2 = -53 a1 + 18d = -58 a1 + 18(-53) = -58 a1 - 954 = -58 a1 = 954 - 58 = 896 So the nth term is: \(\Large a_n = 896 -58(n-1)\)

OpenStudy (aum):

Sorry. The nth term is: \(\Large a_n = 896 -53(n-1)\)

OpenStudy (anonymous):

Thank you so much I totally understand now!!

OpenStudy (aum):

You are welcome.